Martinez Companys relevant range of production is 7,500 units to 12,500 units. When it produces and sells 10,000 units, its average costs per unit are as follows Direct materials Direct labor Variable manufacturing overhead Fixed manufacturing overhead Fixed selling expense Fixed administrative expensee Sales commissions Variable administrative expense Average Cost Per Unit $5.60 $3.10 $1.40 $4.00 $2.60 $2.20 $1.20 $0.45 12. If 12,500 units are produced, what is the total amount of manufacturing overhead cost incurred to support this level of production? What is this total amount expressed on a per unit basis? (Round your per unit answer to 2 decimal places and other answers to the nearest whole dollar amount.) Total manufacturing overhead cost Manufacturing overhead per unit[The following information applies to the questions displayed below.] Martinez Companys relevant range of production is 7,500 units to 12,500 units. When it produces and sells 10,000 units, its average costs per unit are as follows Direct materials Direct labor Variable manufacturing overhead Fixed manufacturing overhead Fixed selling expense Fixed administrative expense Sales commissions Variable administrative expense Average Cost Per Unit $5.60 $3.10 $1.40 $4.00 $2.60 $2.20 $1.20 $0.45 13. If the selling price is $21.60 per unit, what is the contribution margin per unit? (Do not round intermediate calculations. Round your answer to 2 decimal places.) Contribution margin per unitMartinez Companys relevant range of production is 7,500 units to 12,500 units. When it produces and sells 10,000 units, its average costs per unit are as follows Direct materials Direct labor Variable manufacturing overhead Fixed manufacturing overhead Fixed selling expense Fixed administrative expense Sales commissions Variable administrative expense Average Cost Per Unit $5.60 $3.10 $1.40 $4.00 $2.60 $2.20 $1.20 $0.45 14. If 10,000 units are produced, what are the total amounts of direct and indirect manufacturing costs incurred to support this level of production? (Do not round intermediate calculations.) Total direct manufacturing cost Total indirect manufacturing costMartinez Companys relevant range of production is 7,500 units to 12,500 units. When it produces and sells 10,000 units, its average costs per unit are as follows Direct materials Direct labor Variable manufacturing overhead Fixed manufacturing overhead Fixed selling expense Fixed administrative expense Sales commissions Variable administrative expense Average Cost Per Unit $5.60 $3.10 $1.40 $4.00 $2.60 $2.20 $1.20 $0.45 15. What incremental manufacturing cost will Martinez incur if it increases production from 10,000 to 10,001 units? (Round your answer to 2 decimal places.) Incremental cost per unit produced
Answer #1
12
Variable manufacturing overhead 17500 =12500*1.4
Fixed manufacturing overhead 40000 =10000*4
Total manufacturing overhead cost 57500
Manufacturing overhead per unit 4.60 =57500/12500
13
Selling price 21.60
Less: Unit variable cost 11.75 =5.6+3.1+1.4+1.2+0.45
Contribution margin per unit 9.85
14
Direct materials 56000 =10000*5.6
Direct labor 31000 =10000*3.1
Total direct manufacturing cost 87000
Variable manufacturing overhead 14000 =10000*1.4
Fixed manufacturing overhead 40000 =10000*4
Total indirect manufacturing cost 54000
15
Direct materials 5.60
Direct labor 3.10
Variable manufacturing overhead 1.40
Incremental cost per unit produced 10.10